OpenStudy (anonymous):

An express and local train leave Grayslake a 3pm and head for Chicago 50 miles away. The express travels fast as the local and arrives 1 hours ahead of the local. Find the speed of each train...Ihow do i write out two equations for the problem.

OpenStudy (anonymous):

When do they reach Chicago?

OpenStudy (anonymous):

Doesnt say

OpenStudy (anonymous):

\[\left\{\text{express}=\frac{50}{t},\text{local}=\frac{50}{1+t}\right\} \]